# Semantic-guided Search for Efficient Program Repair with Large Language Models

### v1

@@ 4.1. Research Questions
Continuing from our motivation study in Section 2, our empirical evaluation in this section aims to answer the following four research questions:
  • RQ3: How effective is FLAMES? We assess the effectiveness of FLAMES on the Defects4J (Just et al., 2014) and HumanEval-Java (Jiang et al., 2023a) datasets, comparing it to 15 leading APR techniques described in Section 4.2.
  • RQ4: How efficient is FLAMES? We evaluate the time and memory efficiency of FLAMES and compare it to the conventional beam search algorithm.
  • RQ5: How effective is FLAMES’s semantic-guided patch generation? We investigate the effectiveness of FLAMES’s semantic-guided patch generation by implementing our approach with five different LLM-based APR techniques and comparing their effectiveness to that of the original models using two widely-used patch generation strategies: beam search (Ye et al., 2022a; Silva et al., 2023) and multiple sampling (Wei et al., 2023; Xia et al., 2023a).
  • RQ6: How do different configurations affect the performance of FLAMES? Finally, we examine the impact of various key factors on FLAMES’s effectiveness, including reward function, expansion size, and policy algorithm.
@@ 4.2. Experimental Setup
Similar to our initial motivation study (Section 2.1), we evaluate our method using 333 single-hunk bugs from the Defects4J dataset (Just et al., 2014). We also employ the HumanEval-Java dataset, which includes 163 artificial bugs introduced by Jiang et al.(Jiang et al., 2023a). This dataset, recently developed by injecting bugs into the HumanEval benchmark(Chen et al., 2021c), presents a reduced risk of data leakage, where evaluation data might overlap with training data, compared to Defects4J.
@@
To evaluate effectiveness of our approach (RQ3), we compare FLAMES against leading baselines across various APR categories: Template-based, NMT-based, Cloze-based, and LLM-based APR techniques. For LLM-based APR, we evaluate five top models—CodeGen-2B, CodeGen-6B, InCoder-1B, InCoder-6B, and RepairLlama—using their best variants identified in RQ2 with beam sizes from 10 to 200. For Cloze-based APR, we include FitRepair (Xia et al., 2023a), AlphaRepair (Xia and Zhang, 2022), and Repilot (Wei et al., 2023). We also compare FLAMES with NMT-based methods: ITER (Ye and Monperrus, 2024), Recoder (Zhu et al., 2021), KNOD (Jiang et al., 2023b), and RewardRepair (Ye et al., 2022a), as well as template-based approaches: GAMMA (Zhang et al., 2023b), TENURE (Meng et al., 2023), and TBar (Liu et al., 2019). Following standard practices (Jiang et al., 2023a, 2021; Chen et al., 2021a; Wei et al., 2023; Xia and Zhang, 2022), we perform evaluations under the assumption of perfect fault localization, which ensures that variations in fault localization techniques do not affect the results. For LLM-based techniques, we use the same configurations as in Section 3.7. For other techniques, we collect bug fix results from their original papers and replication packages, following common practices in prior works (Jiang et al., 2021; Wei et al., 2023; Xia and Zhang, 2022; Ye et al., 2022a).
@@
To evaluate the effectiveness of FLAMES’s semantic-guided patch generation (RQ5), we also compared our approach two well-known patch generation strategies: beam search and multiple sampling. For beam search, we leverage the standard implementations from Huggingface library,widely adopted toolset for training and deploying LLMs. For multiple sampling, we first set temperature of LLM to 1 for maximizing their diversity. Then, we repeat the default patch generation of these models (using beam search) until reaching 200 patches.
@@
We assess the effectiveness of APR techniques using a standard metric: the number of correct patches. To identify patch correctness, following prior works (Ye et al., 2022a; Le et al., 2019; Xia and Zhang, 2022; Lutellier et al., 2020; Le-Cong et al., 2024), we conduct a manual evaluation to determine if they are either syntactically or semantically equivalent to developer-written patches. Note that, while manual patch correctness assessment is precise, it is resource-intensive and necessitates considerable human effort. Therefore, due to limited human resources, we restrict the use of the number of correct patches as the sole metric for evaluating RQ3, which compares the effectiveness of FLAMES with state-of-the-art baselines. For other research questions, we use the number of plausible patches as a proxy for effectiveness, following prior research (Jiang et al., 2023a; Silva et al., 2023).
@@ 4.3. RQ3: Comparison with Existing Techniques
Table 2 reports the number of correct fixes of our proposed technique, FLAMES, and baseline approaches in the Defects4J and HumanEval-Java dataset.
Table 2. Number of correctly fixed bugs by each tool across Defects4J and HumanEval-Java benchmarks.
Techniques Direction Defects4J HumanEval-Java
TBar Template 46 -
GAMMA Template 87 -
TENURE Template 84 -
RewardRepair NMT 79 -
Recoder NMT 43 -
KNOD NMT 110 -
ITER NMT 55 -
AlphaRepair Cloze 95 -
FitRepair Cloze 123 -
Repilot Cloze 116 -
InCoder-1B LLM 57 82
InCoder-6B LLM 50 71
CodeGen-2B LLM 41 60
CodeGen-6B LLM 49 57
RepairLLama LLM 112 92
FLAMES (Ours) LLM 133 103
4.3.1. Repaired Bugs.
Overall, our approach, FLAMES, successfully generates correct fixes for 133 out of 333 single-hunk bugs in Defects4J, outperforming all existing APR techniques. Notably, it corrects 21 more bugs than RepairLlama, the best-performing LLM-based method, which utilizes the same core LLM as FLAMES. This superior performance is primarily due to the memory limitations of RepairLlama, which frequently crashes due to OOM errors with larger beam sizes, as discussed in Section 2.2. Consequently, it cannot scale to higher beam sizes, with the best-performing model capped at a beam size of 25. This limitation results in a lower number of validated patches of 25. In contrast, FLAMES can generate up to 200 patches, addressing scalability issues and boosting repair success rates.
Moreover, our proposed technique also outperforms the leading cloze-based approach, FitRepair, and the top NMT-based technique, KNOD, by repairing 10 and 23 more bugs respectively, corresponding to improvement rates of 8% and 21%. It is noteworthy that FLAMES achieves this performance with a budget of only 200 validated patches and a 30-minute timeout for each bug, while FitRepair and KNOD require 1000-5000 validated patches and a 5-hour timeout for each bug.
4.3.2. Unique Repaired Bugs.
Next, to further demonstrate the effectiveness of our approach, we evaluate the number of unique bugs that only FLAMES can fix. Figure 3 illustrates our uniqueness analysis on Defects4J dataset of our approach with respects to the best baselines for NMT-based (KNOD (Jiang et al., 2023b)), Cloze-based (FitRepair (Xia et al., 2023a)) and LLM-based (RepairLLama (Silva et al., 2023)) APR approaches.
Figure 3. Unique Repaired Bugs in Defects4J
From Figure 3, we can see that FLAMES can fix 14 unique bugs in Defects4J dataset, which were not fixed by any state-of-the-art baselines. Listing 1 illustrates the unique correct patches generated by FLAMES for the Cli-4 in the Defects4J dataset. These patches involve highly complex additional code, including intricate if-else statements, making them challenging to generate with an insufficient patch space. Consequently, LLM-based approaches like RepairLLama, which are limited by memory constraints, fail to produce correct patches for these bugs.
Furthermore, although cloze-based, NMT-based, and template-based approaches can generate a large number of patches (ranging from 1,000 to 5,000), they still fail to produce the correct ones. We believe this is because these methods rely solely on the models’ output probabilities during patch generation, resulting in inefficiencies when searching for plausible patches. For example, FitRepair generates 5,000 patches and validates the top 1,000 within 5 hours, yet it still fails to produce correct patches for the Cli-4 bug. In contrast, FLAMES incorporates feedback from validation directly into the patch generation process using PG-TD, offering valuable guidance for creating correct patches. As a result, FLAMES can generate patches for these complex bugs with only 200 validated patches and a 30-minute timeout.
@@
Additionally, we also found that FLAMES achieve the highest number of unique fixes among baselines. Particularly, RepairLlama, KNOD, and FitRepair contribute 6, 5, and 9 unique fixes, respectively while FLAMES can fix 14 unique bugs. Moreover, there are 85 correct fixes that are not addressed by these four leading techniques. By integrating all techniques, we successfully fix 299 bugs. This demonstrates that FLAMES can be combined with other methods to significantly increase the number of correct fixes generated.
@@
Finally, to address concerns of data leakage (Jiang et al., 2023a) and benchmark overfitting (Durieux et al., 2019), we extend our evaluation of FLAMES to include HumanEval-Java. Due to the lack of data on existing techniques in the literature, our analysis primarily compares the effectiveness of FLAMES with LLM-based approaches. Detailed results are presented in Table 2.
Listing 1: FLAMES’s correct patch for the bug Cli-4 in Defects4J dataset
if (requiredOptions.size() > 0)
Iterator iter = requiredOptions.iterator();
+ StringBuffer buff = new StringBuffer("Missing required option");
+ if (requiredOptions.size() > 1)
+ buff.append("s");
+ buff.append(": ");
- StringBuffer buff = new StringBuffer();
FLAMES achieves the highest success rate, correctly fixing 103 out of 164 bugs in the HumanEval dataset. Specifically, FLAMES fixes 11 more bugs than the best-performing baseline, RepairLlama, which generates 92 correct fixes, showing an improvement of 12%. Compared to other models, FLAMES is 26% and 45% more effective than InCoder-1B and InCoder-6B, respectively. This performance gap widens when compared to the CodeGen variants; FLAMES outperforms CodeGen-2B and CodeGen-1B by approximately 72% and 81%, respectively.
Answers to RQ3 (Effectiveness): FLAMES successfully fixes 133 and 103 bugs in the Defects4J and HumanEval-Java benchmarks, outperforming current state-of-the-art APR techniques. Additionally, FLAMES uniquely fixes 14 bugs in the Defects4J dataset that cannot be fixed by existing APR techniques.
@@ 4.4. RQ4: Efficiency
Table 4 compares the memory usage of FLAMES with three baselines: (1) the original LLM with beam search, (2) the quantized LLM with beam search, and (3) the original LLM with sequential beam search, across five different LLMs. We evaluate VRAM utilization using three key metrics: average VRAM usage, peak VRAM usage observed in our dataset, and out-of-memory (OOM) ratios. Since our hardware is capped at 80GB VRAM, we assign a memory usage of 80GB to data points that experience OOM errors. Consequently, in cases of OOM, the reported memory usage represent a lower bound of the actual values.
Table 3. Comparison of memory usage among the original LLM (Original), quantized LLM (Quantized), original LLM with sequential beam search (SeqBS) and FLAMES.
@@
Flames 32.4 37.4 0.0
Original 66.6 80 62.8
RepairLlama Quantized - - -
SeqBS - - -
Flames 8.1 22.8 0.0
Table 4. Comparison of average running times of FLAMES and beam search (BS) for finding plausible patches with statistical measures on different models and beam sizes
Model Beam Avg. Time (s) $|\mathbf{\delta}|$ p-value
@@
We can see that the average VRAM usage of FLAMES shows a substantial decrease across all models evaluated. For instance, with the CodeGen-2B model, average usage drops from at least 72.2 GB in the original configuration to just 12.1 GB using FLAMES with a reduction of over 83%. This trend of reduced VRAM usage is consistent across other models, such as the Incoder-1B, where average usage is reduced from at least 52.7 GB to 7.4 GB.
In terms of peak VRAM usage, while the baselines frequently reach the 80 GB VRAM limit, FLAMES maintains peak values significantly well below this threshold, such as 13.9 GB for CodeGen-2B and 37.4 GB for Incoder-6B. This effective management of peak VRAM not only avoids bottlenecks and crashes from OOM errors but also enhances the practical usability of these models on limited-resource hardware environments.
As a result, the FLAMES method achieves a zero out-of-memory (OOM) ratio across all models, in contrast to the higher OOM ratios observed with the original configurations. For example, the CodeGen-6B model under the original methodology reports an OOM ratio of 80.5%, which significantly limits model reliability. The absence of OOM errors with FLAMES suggests a notable enhancement in the stability and memory efficiency of LLMs.
Answers to RQ4.1 (Memory Efficiency): FLAMES can significantly reduce VRAM usage in LLMs, successfully minimizing average VRAM usage by up to 83% and reducing peak VRAM requirements from 80 GB to as low as 12.7 GB across various configurations and models. Additionally, FLAMES do not introduce any OOM errors while baselines witness OOMs in up to 86.6% of data points.
@@
In this experiment, we evaluate the time efficiency of LLM-based APR using our proposed method, FLAMES, compared to traditional beam search techniques. Our aim is to validate the hypothesis that ”FLAMES can identify plausible patches more quickly than the traditional beam search algorithm”. To validate this hypothesis, we measure the time each method takes to successfully identify plausible patches, employing the Mann-Whitney-Wilcoxon (MWW) statistical test and Cliff’s Delta for effect size assessment. This experiment focuses on the CodeGen-2B, InCoder-1B, and RepairLlama models, chosen for their lower incidence of out-of-memory (OOM) crashes, ensuring a robust sample size for reliable testing. The detailed results are presented in Table 4.
For the CodeGen-2B model, our hypothesis is substantiated across all beam sizes, with p-values ranging from 0.002 to 0.013 and effect sizes between 0.22 and 0.37, indicating that FLAMES is consistently more time-efficient than beam search. Specifically, FLAMES shows lower running times across all beam sizes, reducing time by approximately 11.44 seconds at a beam size of 10 and up to nearly 39.52 seconds at a beam size of 200.
The RepairLlama model also supports the hypothesis with p-values between 0.001 and 0.027 across all beam sizes, confirming the time efficiency of FLAMES over beam search, although the effect sizes range from negligible to small. Consequently, FLAMES demonstrates lower average running times at the beam sizes of 10 and 200. For other beam sizes, while FLAMES generally performs faster than beam search, occasional longer run times by FLAMES skew the average, leading to variability in results.
In contrast, the InCoder-1B model exhibits greater variability between the two methods. While FLAMES tends to generate patches faster at higher beam sizes (50, 100, and 200), it underperforms at a beam size of 25 and shows no significant difference at 200. The most favorable performance for FLAMES is observed at a beam size of 10, where it significantly outperforms beam search (p-value of 0.001). However, for beam sizes 25 and 200, the differences in running times lack statistical significance, indicating that the benefits of FLAMES do not uniformly apply across all configurations within this model.
Answers to RQ4.2 (Time Efficiency): FLAMES significantly improves time efficiency in LLMs, surpassing traditional beam search methods.
4.5. RQ5: Effectiveness of semantic-guided patch generation

### v2

@@ 4.1. Research Questions
Continuing from our motivation study in Section 2, our empirical evaluation aims to answer the following four research questions.
RQ3: How effective is FLAMES? We assess the effectiveness of FLAMES on the Defects4J (Just et al., 2014), HumanEval-Java (Jiang et al., 2023a) and TransformedD4J (Le-Cong et al., 2025) datasets, comparing our proposed technique to 15 state-of-the-art APR techniques and Qwen2.5-Coder-32B, the best open-source LLM for code in well-known leaderboards (Zhuo et al., 2024; Liu et al., 2023).
RQ4: How efficient is FLAMES? We evaluate the time and memory efficiency of FLAMES and compare it with standard and sequential beam search algorithms.
RQ5: How effective is FLAMES’s semantic-guided patch generation? We investigate the effectiveness of FLAMES’s semantic-guided patch generation by implementing our approach with six LLMs for APR and comparing their effectiveness to the full-precision models using two widely used patch generation strategies: beam search (Ye et al., 2022a; Silva et al., 2023) and multiple sampling (Wei et al., 2023; Xia et al., 2023a).
RQ6: How do different configurations affect the performance of FLAMES? Finally, we examine the impact of various key factors on the effectiveness of FLAMES, including the reward function, the expansion size, and the policy algorithm.
@@ 4.2. Experimental Setup
Similar to our initial motivation study (Section 2.1), we evaluated our method using 333 single-hunk bugs from the Defects4J dataset (Just et al., 2014). Furthermore, we used HumanEval-Java (Jiang et al., 2023a) and TransformedD4J (Le-Cong et al., 2025), which contain artificial bugs introduced by applying bug injection to normal programs and semantic-preserving transformations to Defects4J bugs, respectively, to mitigate the risk of data leakage.
@@
To evaluate the effectiveness of our approach (RQ3), we compared FLAMES with the leading baselines in various APR categories: Template-based, NMT-based, Cloze-based, and LLM-based APR techniques. For LLM-based APR, we evaluated five models using their best variants identified in RQ2. For Cloze-based APR, we included FitRepair (Xia et al., 2023a), AlphaRepair (Xia and Zhang, 2022), and Repilot (Wei et al., 2023). We also compared FLAMES with NMT-based methods, i.e., ITER (Ye and Monperrus, 2024), Recoder (Zhu et al., 2021), KNOD (Jiang et al., 2023b), and RewardRepair (Ye et al., 2022a), as well as template-based approaches, i.e., GAMMA (Zhang et al., 2023b), TENURE (Meng et al., 2023), and TBar (Liu et al., 2019). Finally, we also included Qwen2.5-Coder-32B (Hui et al., 2024), the leading LLM for code on well-known leaderboards (Liu et al., 2023; Zhuo et al., 2024). Following standard practices (Jiang et al., 2023a, 2021; Chen et al., 2021a; Wei et al., 2023; Xia and Zhang, 2022), we performed evaluations under the assumption of perfect fault localization, which ensured that variations in fault localization techniques do not affect the results. For LLM-based techniques, we used the same configurations as in Section 3.6. For other techniques, we collected bug fix results from their original papers and replication packages, following common practices in prior works (Jiang et al., 2021; Wei et al., 2023; Xia and Zhang, 2022; Ye et al., 2022a).
@@
To evaluate the effectiveness of FLAMES’s semantic-guided patch generation (RQ5), we also compared our approach with two well-known patch generation strategies: beam search and multiple sampling. For beam search, we leverage the standard implementations from the Huggingface library, a widely adopted toolset for training and deploying LLMs. For multiple sampling, we first set the LLM temperature to 1 to maximize its diversity. Then, we repeat the default patch generation of these models (using beam search) until we reach 200 patches.
@@
We assessed the effectiveness of APR techniques using a standard metric: the number of correct patches. To identify patch correctness, following prior works (Ye et al., 2022a; Le et al., 2019; Xia and Zhang, 2022; Lutellier et al., 2020; Le-Cong et al., 2025), we conducted a manual evaluation to determine if they were syntactically or semantically equivalent to developer-written patches. Note that while manual patch correctness assessment is precise, it is resource-intensive and requires considerable human effort.
@@ 4.3. RQ3: Comparison with Existing Techniques
Table 2. Comparison with state-of-the-art APR techniques on Defects4J benchmark. #Correct and #Plausible denotes the number of correct and plausible patches generated by each tool. RepairLLama (+BS) and (+MS) denotes the results of RepairLLama with beam search and multiple sampling, respectively.
Techniques #Correct #Plausible Precision
TBar 46 - -
GAMMA 87 - -
TENURE 84 - -
RewardRepair 79 - -
Recoder 43 - -
KNOD 110 - -
ITER 55 - -
AlphaRepair 95 - -
FitRepair 123 - -
Repilot 116 - -
InCoder-1B 42 75 0.56
InCoder-6B 50 96 0.52
CodeGen-2B 41 88 0.47
CodeGen-6B 49 94 0.52
Qwen2.5-Coder-32B 68 112 0.61
RepairLLama (+BS) 112 170 0.66
RepairLLama (+MS) 90 151 0.59
FLAMES 133 211 0.63
4.3.1. Repair Success.
Table 2 presents the number of correct fixes generated by our approach, FLAMES, and baseline methods in Defects4J. FLAMES successfully repairs 133 of the 333 single-hunk bugs, outperforming all existing APR techniques.
Comparison with RepairLLama’s Variants. Specifically, our approach generates correct patches for 21 and 43 more bugs than two other variants with beam search and multiple sampling of its core LLM, RepairLLama. These results highlight the advantages of our semantic-guided patch generation approach using PG-TD. This method enables FLAMES to scale to larger beam sizes without compromising memory efficiency, as observed in RepairLLama with beam search. Meanwhile, it also guides the patch generation process with semantic feedback, resulting in a more effective approach than multiple sampling without guidance. Although this approach slightly reduces the precision from 0.66 to 0.63, this trade-off is justified by a 19% increase in repair success. Moreover, FLAMES maintains a higher precision than other LLM-based APR techniques and RepairLLama with multiple sampling.
Comparison with State-of-the-Art Techniques. Additionally, FLAMES significantly outperforms other state-of-the-art APR baselines. It correctly fixes 10 more bugs than FitRepair, the best baseline, achieving an 8% improvement. Additionally, FLAMES outperforms the top NMT-based technique, KNOD, and the top template-based approach, GAMMA, by 21% and 53%, respectively, in the number of correct patches. Notably, FLAMES achieves this performance with a budget of only 200 validated patches and a 30-minute pause per bug, while FitRepair and KNOD require 1,000–5,000 validated patches and a pause of 5-hours per bug. Furthermore, FLAMES outperforms Qwen2.5-Coder-32B, the leading LLM for code, by 51%. This advantage comes from two key factors: first, this model is not specifically trained for APR, unlike our core model, RepairLlama; second, FLAMES incorporates semantic-guided patch generation that enhances its repair effectiveness.
@@
Unique Fixes. We found that FLAMES achieves the highest number of unique fixes among the best baselines for NMT-based (KNOD (Jiang et al., 2023b)), Cloze-based (FitRepair (Xia et al., 2023a)) and LLM-based (RepairLLama (Silva et al., 2023)) APR approaches. Particularly, RepairLlama, KNOD, and FitRepair contribute 6, 5, and 9 unique fixes, respectively, while FLAMES can fix 14 unique bugs. Moreover, there are 85 correct fixes that are not addressed by these four leading techniques. By combining all the techniques, we can successfully fix 299 bugs. This shows that FLAMES can be complementarily used with existing APR methods such as KNOD and FitRepair to significantly increase the number of correct fixes generated.
Table 3. Comparison with RepairLLama on multi-location bugs from Defects4J.
Techniques Correct/Plausible Patches
RepairLLama 16/35
FLAMES 24/59
4.3.2. Repair Success on Multi-Location Bugs.
Beyond the comparison on single-location bugs, we also conduct experiments on multi-location bugs following the settings of RepairLLama (Silva et al., 2023) with 172 bugs from Defects4J. In this setting, each bug requires edits on multiple non-consecutive locations within a single function. We compare the effectiveness of FLAMES with our base model and also the best LLM baseline, RepairLlama to further understand the impact of FLAMES on repairing multi-location bugs. The detailed results are presented in Table 3.
Our experimental results show that while RepairLLama (the best LLM baseline) generates 35 plausible patches for these bugs, FLAMES (with RepairLLama as the base LLM) can generate 59 plausible patches, demonstrating 68% improvement over the best baseline. Further inspection reveals that FLAMES generates 24 correct patches out of 59 plausible ones, compared to only 16 correct patches generated by RepairLLama. This result confirms the advantages of our patch generation for such complex scenarios.
Table 4. Comparison with state-of-the-art APR techniques on HumanEval-Java and TransformedD4J benchmark. The results are displayed as x/y, with x, y denotes the number of bugs with correct and plausible patches, respectively.
Techniques HumanEval-Java TransformedD4J
CodeGen-6B 57/65 _/393
InCoder-6B 71/81 _/427
Qwen2.5-Coder-32B 68/111 _/403
RepairLLama 92/118 334/581
FLAMES 103/133 456/776
@@
To address concerns related to data leakage (Jiang et al., 2023a) and benchmark overfitting (Durieux et al., 2019), we extended our evaluation of FLAMES and the four most effective LLM-based APR techniques from the previous experiment. Specifically, we evaluated their performance on the HumanEval-Java (Jiang et al., 2021) and TransformedD4J (Le-Cong et al., 2025) benchmarks. Given the large number of plausible patches generated by these techniques on TransformedD4J, we limited our correctness assessment to patches produced by the two most effective methods, namely FLAMES and RepairLLama.
Overall, FLAMES significantly outperforms the baselines in terms of the number of correct patches. For example, in HumanEval-Java, FLAMES successfully fixes 11 more bugs than the best-performing baseline, RepairLlama, which generates 92 correct fixes, representing an improvement of 12%. Similarly, in TransformedD4J, FLAMES achieves a 36.5% improvement over the best baseline. These results demonstrate the generalizability of FLAMES across various evaluation benchmarks, highlighting its effectiveness in generating correct fixes in different data sets.
Answers to RQ3 (Effectiveness): FLAMES successfully fixes 133, 103, and 456 bugs on Defects4J, HumanEval-Java, and TransformedD4J, significantly outperforming the best baselines by 8%, 12% and 36. 5%, respectively.
@@ 4.4. RQ4: Efficiency
Table 5 compares the memory usage of FLAMES with three baselines: (1) the full-precision LLM with beam search, (2) the quantized LLM with beam search, and (3) the full-precision LLM with sequential beam search, across five different LLMs. We evaluated VRAM usage using three key metrics: average VRAM usage, peak VRAM usage observed in our dataset, and out-of-memory ratios. Since our hardware is capped at 80GB VRAM, we assign a memory usage of 80GB to data points that experience OOM errors. Consequently, in cases of OOM, the reported memory usage represents a lower bound of the actual values.
Table 5. Comparison of memory usage among the full-precision LLM (Full-Precision), quantized LLM (Quantized), full-precision LLM with sequential beam search (SeqBS) and full-precision LLM with FLAMES.
@@
PG-TD 32.4 37.4 0.0
Full-Precision 66.6 80 62.8
RepairLlama PG-TD 8.1 22.8 0.0
Overall, the VRAM usage of FLAMES demonstrates a substantial reduction across all evaluated models, outperforming both the original full-precision configurations and the two widely used memory reduction techniques: Quantization and Sequential Beam Search (SeqBS). For example, in the case of the InCoder-1B model, the average VRAM usage drops from 52.7 GB in the original configuration to just 7.4 GB with FLAMES, representing a reduction of 86%. In comparison, the quantization approach reduces memory usage to 31.2 GB, while SeqBS increases it to 56.1 GB, making FLAMES the most memory-efficient method.
Regarding peak VRAM usage, while baseline methods frequently reach the 80 GB VRAM limit, FLAMES maintains peak values well below this threshold. Notably, FLAMES achieves a maximum VRAM usage of just 12.7 GB for the smallest model (InCoder-1B) and 37.4 GB for the largest model (InCoder-6B). In contrast, both the original full-precision and SeqBS configurations consistently hit the 80 GB VRAM limits, increasing the risk of out-of-memory (OOM) errors.
Furthermore, in terms of OOM ratios, FLAMES completely eliminates OOM errors across all models, achieving a 0% OOM ratio, while the original full-precision and SeqBS methods frequently experience severe memory failures. For example, the full-precision CodeGen-6B model encounters an 80.5% OOM rate, making it unusable in memory-constrained environments. Even quantization, while reducing memory usage, does not guarantee the elimination of OOM errors, as observed with CodeGen-6B, which still reports an OOM ratio of 49.8%.
Answers to RQ4.1 (Memory Efficiency): FLAMES can significantly reduce VRAM usage in LLMs, successfully minimizing average VRAM usage by up to 83% and reducing peak VRAM requirements from 80 GB to as low as 12.7 GB across various configurations and models. Additionally, FLAMES eliminates OOM errors, while the baselines witness OOMs in up to 86.6% of the data points.
Table 6. Comparison of average running times of FLAMES and beam search (BS) for finding plausible patches with statistical measures on different models and number of generated patches (# Patches)
Model # Patches Avg. Time (s) $|\mathbf{\delta}|$ p-value
@@
In this experiment, we evaluated the time efficiency of LLM-based APR using our proposed method, FLAMES, compared to traditional beam search techniques. Our aim is to validate the hypothesis that ”FLAMES can identify plausible patches more quickly than the traditional beam search algorithm”. To validate this hypothesis, we measure the time each method takes to successfully identify plausible patches, employing the Mann-Whitney-Wilcoxon (MWW) statistical test and Cliff’s Delta for effect size assessment. This experiment focuses on the CodeGen-2B, InCoder-1B, and RepairLlama models, chosen for their lower incidence of out-of-memory (OOM) crashes, ensuring a robust sample size for statistical testing. The detailed results are presented in Table 6.
For the CodeGen-2B model, our hypothesis is supported in all examined numbers of generated patches, with p values between 0.002 and 0.013 and effect sizes between 0.22 and 0.37, indicating that FLAMES is more time efficient than beam search. Notably, FLAMES reduce time by approximately 11.44 seconds and by 39.52 seconds with 10 and 200 generated patches.
The RepairLlama model also supports the hypothesis with p-values between 0.001 and 0.027 across all examined numbers of generated patches, confirming the time efficiency of FLAMES over the beam search, although the effect sizes range from negligible to small. Consequently, FLAMES shows lower average running times with 10 and 200 generated patches. For other configurations, while FLAMES generally performs faster than the beam search, the occasional longer run times of FLAMES skew the average, leading to variability in the results.
In contrast, the InCoder-1B model exhibits greater variability between the two methods. Although FLAMES tends to generate patches faster at higher numbers of generated patches (50, 100, and 200), it performs worse than BS with 25 generated patches. The most favorable performance for FLAMES is observed with 10 generated patches, where it significantly outperforms the beam search (p-value 0.001). However, for 25 and 200 generated patches, the differences in running times lack statistical significance, indicating that the benefits of FLAMES do not apply uniformly in all configurations within this model.
Answers to RQ4.2 (Time Efficiency): FLAMES significantly improve time efficiency in LLMs, surpassing traditional beam search methods.
4.5. RQ5: Effectiveness of Semantic Guided Patch Generation
In this experiment, we evaluated the effectiveness of FLAMES’s semantic-guided patch generation in comparison to beam search and multiple sampling. These methods were evaluated in five LLM-based automated program repair (APR) models, as introduced in Section 2.1.2, together with Qwen2.5-Coder-32B, a state-of-the-art LLM for code. The detailed results are presented in Table 7.
FLAMES consistently outperforms beam search in all models, with improvement rates ranging from 12% to 48% in terms of the number of correct patches. For example, FLAMES generates 21 more correct patches than RepairLlama, our default base model. This is due to FLAMES ’s effective use of GPU resources which allows for a more thorough exploration of the large patch space. In contrast, beam search requires substantial VRAM, resulting in a high out-of-memory crash rate, as discussed in RQ4 and RQ1.